# Practical Collaborative Perception: A Framework for Asynchronous and Multi-Agent 3D Object Detection

### v1

@@ IV-B1 Dataset and Metric
To evaluate our collaboration framework, we use the V2X-Sim 2.0 [18] which is made using CARLA and the traffic simulator SUMO. This dataset is made of 100 100-frame sequences of traffic taking place at intersections of three towns of CARLA which are Town 3, Town 4, and Town 5. Each sequence contains data samples recorded at 5 Hz. Each data sample comprises raw sensory measurements made by the ego vehicle, one to four CAVs, and an IRSU, which is placed at an elevated position that has a large minimally occluded field of view of the intersection. Every vehicle and the IRSU are equipped with a 32-channel LiDAR. All agents are in sync which results in the same timestamp of data that they collect.
@@ IV-B2 Implementation, Experiments and Results
In the convention of the V2X-Sim dataset, the IRSU and the ego vehicle are respectively assigned the id of 0 and 1 while other CAVs get id ranging from 2 to 5. To test our collaboration method’s ability to handle asynchronous exchanged information, we set the time lag between the timestamp $t$ of the ego vehicle’s query and the timestamp $t_{i}$ of the detection $\mathcal{B}_{i}=\left\{\mathbf{b}_{i,j}\right\}$ made by agent $\mathcal{A}_{i}(i\in\{0,2,3,4,5\})$ to the time gap between two consecutive data sample of V2X-Sim which is $0.2$ seconds.
@@
On the other hand, Early Collaboration reaches high performance by exchanging the entire raw point cloud sequences $\mathcal{S}_{i}=\left\{\mathcal{P}^{t_{i}-K+1}_{i},\dots,\mathcal{P}^{t_{i}}_{i}\right\}$ collected by each agent $\mathcal{A}_{i}$ is the second baseline. To explore the upper bound of V2X collaborative perception, we only apply the synchronous exchange to this baseline.
To verify the ability of enhancing single-vehicle perception of V2X collaboration, we evaluate our approach and baselines in the setting where ground truths are made of objects visible to the ego vehicle meaning their bounding boxes contain at least one point of the point cloud $\mathcal{P}_{e}^{t}$, which the ego vehicle obtained at the time step of query $t$. The result of this evaluation is summarized in Tab.V which shows that late collaboration in all three settings is not beneficial, as the best late collaboration is 4.42 mAP behind the single-agent perception. This is because True Positive (TP) detections made by other agents are counted as false positives if they are not visible to the ego vehicle. In addition, ill-localized but overly confident detections made by other agents can suppress good detections made by the ego vehicle, thus further reducing the overall performance. Remarkably, our collaboration approach using MoDAR points outperforms the early collaboration by 3.35 mAP while consuming the same amount of bandwidth as late collaboration. This reaffirms our design philosophy that a good mutli-agent collaborative perception framework can be made on the foundation of a good single-agent perception model and a simple collaboration method.
@@
Collaboration Method mAP Exchanged Info
Size (MByte)
None 65.71 0
Late - async 54.77 $0.01$
Late - async prop. 59.34 $0.01$
Late - sync 61.29 $0.01$
Mid - DiscoNet 69.98 -
Early 71.84 $33.95$
Ours - MoDAR 75.19 $0.01$
Beside enabling the ego vehicle more accurately detecting objects that are visible to itself, another great benefit of collaborative perception is to help the ego vehicle see the invisible objects - those do not contain any points of its point cloud $\mathcal{P}_{e}^{t}$, thus overcoming the challenge of occlusion and sparsity. To demonstrate this, we perform the relax the criterion of eligibility of ground truth such that they only need to contain at least one point emitted by any agents in the V2X network. The performance shown in Tab.VI indicates that on a larger set of ground truths, Late Collaboration does improve performance, compared to single-vehicle (no collaboration). The improvement is significant (at least $8.35$ mAP) even in the poorest setting where the set of exchanged detection $\left\{\mathcal{B}_{i}|i\in\{0,2,3,4,5\}\right\}$ is $0.2$ seconds behind the time of query, thus resulting a spatial misalignment between detected objects that are exchanged and their associated ground truths if the underlying objects are dynamic. This can be explained by a significant number of static and slow-moving objects present in an intersection whose past detections remain true positive at the present. When detections of dynamic objects are better accounted for as in the sync setting where agents exchange their detections at the same timestamp as the query, the performance is largely improved by $9.29$ mAP, compared to the async setting. However, this sync setting is unrealistic because different agents have different detection rates. Interestingly, propagating detections using scene flow as in the async prop. setting can reach $96.20\%$ the performance of the sync setting ($2.68$ mAP behind). This implies the effectiveness of scene flow estimated by our Aligner++. In this setting, our collaboration method using MoDAR no longer outperfroms the early collaboration baseline as some distant objects (with respect to the ego vehicle) are inevitably missed by connected agents, thus remaining invisible to the ego vehicles and being unable to be detected. However, our method still maintain favorable performance of $98.2\%$ of the early collaboration’s while attaining as low bandwidth usage as late collaboration.
TABLE VI: performance of collaboration methods on ground truth that are visible to at least one agents
@@
None 52.84
Late - async 61.19
Late - async prop. 67.80
Late - sync 70.48
Early 78.10
Ours - MoDAR 76.72
Ours - MoDAR 75.41
and Foreground

### v3

@@ IV-B1 Dataset and Metric
To evaluate our collaboration framework, we use the V2X-Sim 2.0 [19] which is made using CARLA [36] and the traffic simulator SUMO [37]. This dataset is made of 100 100-frame sequences of traffic taking place at intersections of three towns of CARLA which are Town 3, Town 4, and Town 5. Each sequence contains data samples recorded at 5 Hz. Each data sample comprises raw sensory measurements made by the ego vehicle, one to four CAVs, and an IRSU, which is placed at an elevated position that has a large minimally occluded field of view of the intersection. Every vehicle and the IRSU are equipped with a 32-channel LiDAR. All agents are in sync which results in the same timestamp of data that they collect.
@@ IV-B2 Implementation, Experiments and Results
In the convention of the V2X-Sim dataset, the IRSU and the ego vehicle are respectively assigned the identity of 0 and 1 while other CAVs get identities ranging from 2 to 5. To test our collaboration method’s ability to handle asynchronous exchanged information, we set the time lag between the timestamp $t$ of the ego vehicle’s query and the timestamp $t_{i}$ of the detection $\mathcal{B}_{i}=\left\{\mathbf{b}_{i,j}\right\}$ made by agent $\mathcal{A}_{i}(i\in\{0,2,3,4,5\})$ to the time gap between two consecutive data sample of V2X-Sim which is $0.2$ seconds.
Enhance Detection of Visible Objects
@@
On the other hand, Early Collaboration reaches high performance by exchanging the entire raw point cloud sequences $\mathcal{S}_{i}=\left\{\mathcal{P}^{t_{i}-K+1}_{i},\dots,\mathcal{P}^{t_{i}}_{i}\right\}$ collected by each agent $\mathcal{A}_{i}$ is the second baseline.
@@
None 65.71
Late - async 54.77
Late - async prop. 59.34
Late - sync 61.29
Early 71.84
Ours 75.19
To verify the ability to enhance the single-vehicle perception of V2X collaboration, we evaluate our approach and baselines in the setting where ground truths are made of objects visible to the ego vehicle. This means eligible ground truth must contain at least one point of the point cloud $\mathcal{P}_{e}^{t}$, which the ego vehicle obtained at the time step of query $t$. The result of this evaluation is summarized in Tab.V which shows that Late Collaboration in all three settings is not beneficial, as the best Late Collaboration is only $93.3\%$ (4.42 mAP behind) of the single-agent perception. This is because True Positive (TP) detections made by other agents are counted as false positives if they are not visible to the ego vehicle. In addition, ill-localized but overly confident detections made by other agents can suppress good detections made by the ego vehicle, thus further reducing the overall performance. Remarkably, our collaboration approach using MoDAR points outperforms the early collaboration by $4.7\%$ (3.35 mAP). This reaffirms our design philosophy that a good multi-agent collaborative perception framework can be made on the foundation of a good single-agent perception model and a simple collaboration method.
Enhance Detection of Completely Obstructed Regions Objects
Besides enabling the ego vehicle more accurately detect objects that are visible to itself, another great benefit of collaborative perception is to help the ego vehicle see the objects that are completely obstructed - those that do not contain any points of its point cloud $\mathcal{P}_{e}^{t}$, thus overcoming the challenge of occlusion and sparsity. To demonstrate this, we relax the criterion of eligibility of ground truth such that they only need to contain at least one point emitted by any agents in the V2X network. In this experiment, we add a strong mid-collaboration method - DiscoNet [3] to the set of baselines to showcase the capability of our approach compared to the state-of-the-art. In the framework of DiscoNet, connected agents, which have identical detection architecture, exchange BEV images of their point cloud. The ego vehicle fuses the BEV image of its point cloud with the exchanged BEV via a fully connected collaboration graph to use as the input to the detection head.
In addition to the detection precision, we also measure the bandwidth consumption of each collaboration method. Our measurement is based on the raw uncompressed form of the exchanged data to facilitate a fair comparison.
The results shown in Tab.VI indicate that on a larger set of ground truths, Late Collaboration does improve performance, compared to single-vehicle (no collaboration). The improvement is significant (at least $8.35$ mAP) even in the poorest setting where the set of exchanged detection $\left\{\mathcal{B}_{i}|i\in\{0,2,3,4,5\}\right\}$ is $0.2$ seconds behind the time of query, thus resulting a spatial misalignment between detected objects that are exchanged and their associated ground truths if the underlying objects are dynamic. This can be explained by a significant number of static and slow-moving objects present in intersections whose past detections remain true positives at the present. When detections of dynamic objects are perfectly accounted for as in the sync setting where agents exchange their detections at the same timestamp as the query, the performance is largely improved by $9.29$ mAP ($15.2\%$), compared to the async setting. However, this sync setting is unrealistic because different agents have different detection rates. Interestingly, propagating detections using scene flow as in the async prop. setting can reach $96.2\%$ the performance of the sync setting ($2.68$ mAP behind). This implies the effectiveness of scene flow estimated by our Aligner++.
TABLE VI: performance of collaboration methods, measured by mAP, on ground truth that are visible to at least one agent
Collaboration Sync Async Bandwidth
Method Usage (MB)
None 52.84 - 0
Late 70.48 61.19 0.01
67.80  [^1] Exchanged detections are propagated using scene flow according to (5) before fusion by Non-Max Suppression.
Mid - DiscoNet  [^2] Our implementation based on [3] 78.70 73.10 25.16
Early 78.10 77.30 33.95
Ours 79.20 76.72 0.01
In the sync setting, DiscoNet narrowly exceeds Early Collaboration by $0.6$ mAP ($0.7\%$), thus showing the benefit of the deep interaction among agents powered by their dense collaboration graph. However, this benefit comes at the cost of relatively high bandwidth usage of $25$ MB per broadcast BEV image on average which is only $25\%$ less than the average bandwidth required by Early Collaboration. The operation of our method in this setting is as follows: MoDAR points broadcast by other agents have the same timestamp as the point cloud of the ego vehicle; therefore their propagation using scene flow is skipped. Our method slightly exceeds DiscoNet by $0.5$ mAP ($0.6\%$) to reach the highest performance while consuming the least bandwidth among collaborative methods. This confirms the effectiveness of MoDAR in facilitating collaboration among connected agents.
In the async setting, DiscoNet is severely affected by the misalignment among exchanged BEV images caused by timestamp mismatch as its performance dropped by $5.6$ mAP ($7.1\%$). Our method is affected because the misalignment between exchanged MoDAR points and their corresponding ground truth is explicitly taken into account by the propagation of MoDAR points using scene flow. As a result, the performance gap between ours and DiscoNet is extended to $3.62$ mAP ($5\%$). Early Collaboration shows the most robustness against timestamp mismatch with an $0.8$ mAP ($1\%$) drop. This can be explained by the detector’s ability to implicitly handle to some extent the misalignment among point clouds. This phenomenon is well known in the literature on multi-frame single-vehicle object detection [6, 38, 8, 39] where models trained on the concatenation of 0.5-second point cloud sequences enjoy a $30\%$ mAP gain compared to models trained on individual point clouds.
Support Heterogeneous Detector Networks
To show that our collaboration method supports heterogeneous networks out of the box, we perform an experiment where we start with a network of agents uniformly using PointPillar as their detectors and gradually replace PointPillar with SECOND in one agent after another. The result of this experiment is shown in Fig.10.
Figure 10: Performance of the ego vehicle in heterogeneous collaboration networks made PointPillar and SECOND detectors. As the number of SECONDs increases from 0 to 6 - the size of the network, the ego vehicle’s performance transits from its value in a full-PointPillar network to its value in a full-SECOND network without catastrophe dropping out of the lower bound of Late Collaboration.
As the number of SECOND increases, the performance of the collaborative perception gradually transits in a downward trend from $76.72$ mAP of a full-PointPillar network to $71.94$ mAP of a full-SECOND network, without dropping below the lower bound of Late Collaboration, which is the case of mid-collaboration method without dedicated modules for bridging the domain gap [5]. This downward trend is because an individual PointPillar achieves higher AP than an individual SECOND on the V2X-Sim dataset when being trained under the same setting. More interestingly, the combination made of the RSU using SECOND and all connected vehicles using PointPillar archives the best performance of $77.74$ mAP, reaching $99.5\%$ performance of Early Collaboration in sync setting.
Relation between Performance and Network Size
Finally, to study how the performance of collaboration perception evolves with respect to the number of agents in the V2X networks, we gradually increase the number of participants in the collaboration. As can be seen in Fig.11, the collaboration with only the IRSU brings $15.09$ mAP ($28.6\%$) improvement. The magnitude of performance gain reduces and eventually saturated as the number of agents increases. An increased number of agents results in higher overlap in agents’ field of view which makes objects in the overlapped regions well detected by many agents. However, this overlap also means the field of view of the collaboration network as a whole does not increase proportionally with the number of agents. Moreover, exchanging detected objects made by distant agents may not get used by the ego vehicle because they are outside the predefined detection range.
Qualitative
The qualitative performance in Fig.12 shows that collaborative perception using MoDAR points enables the ego vehicle to detect objects that are occluded or have a few to zero LiDAR points due to long range. Particularly, the vehicle at the top of Fig.12a is occluded with respect to the ego vehicle due to the presence of a large vehicle. However, this occluded vehicle is successfully detected thanks to a single MoDAR point produced by the IRSU. Another example of occluded vehicles that are successfully detected is Fig.12e where vehicles in the bottom right corner are occluded by a building-like structure. The illustration of successful detection at long range thanks to MoDAR points can be found in Fig.12b, Fig.12d, and Fig.12f where vehicles near their edges have a few or even zero LiDAR points of the ego vehicle.
Figure 11: The performance of ego vehicle’s detection increases sharply in collaboration with only IRSU, then slows down and saturates at $98.2\%$ of Early Collaboration in sync setting as the number of connected agents increases.
(a) 1 RSU & 3 CAVs
(b) 1 RSU & 1 CAVs
(c) 1 RSU & 3 CAVs
(d) 1 RSU & 2 CAVs
(e) 1 RSU & 3 CAVs
(f) 1 RSU & 1 CAVs
Figure 12: Qualitative performance of our method on three samples of the V2X-Sim dataset. The number and type of connected agents are annotated in the title of each image. Blue points are LiDAR points collected by the ego vehicle. Gray points are LiDAR points collected by other agents which are displayed for visualization purposes only. Orange stars denote the MoDAR points broadcast by other connected agents. Green solid and dashed rectangles respectively represent ground truth visible and invisible to the ego vehicle. Red rectangles are the detections made by the ego vehicle using our method.
V Conclusions and Future Work
In this paper, we develop a collaborative perception framework based on the foundation made of good single-agent perception models and an effective late-early collaboration method. The simplicity of our collaboration method facilitates its practicality by (i) minimizing bandwidth usage, (ii) dismissing the need for inter-agent synchronization, (iii) making minimal changes to single-agent object detectors, and (iv) supporting networks of heterogeneous detectors. Importantly, this practicality does not compromise the overall performance as our collaboration method exceeds Early Collaboration on the V2X-Sim dataset. The success here paves the way for our next development which is the demonstration of our method on a real-world V2X network.
VI Acknowledgement
This work was granted access to the HPC resources of IDRIS under the allocation 2021-AD011012128R1 made by GENCI. This work has been supported in part by the ANR AIby4 under the number ANR-20-THIA-0011. This work was carried out in the framework of the NExT Senior Talent Chair DeepCoSLAM, which is funded by the French Government, through the program Investments for the Future managed by the National Agency for Research (ANR-16-IDEX-0007), and with the support of Région Pays de la Loire and Nantes Métropole.